The sum of a complement and a supplement angle is 200. what is the measure of the angle?
Answer by ankor@dixie-net.com(22740)

The sum of a complement and a supplement angle is 200.
what is the measure of the angle?
(90-a) + (180-a) = 200
-2a = 200 - 270
-2a = 070
a = +35